# Membership Inference for Contrastive Pre-training Models with Text-only PII Queries

### 15 Mar 2026 v1

@@ 4.3 Ablation Study
5 Robustness against Potential Defenses
@@
Table 3: UMID robustness evaluation under differential privacy defense ($\epsilon=1.0$, $\delta=10^{-5}$).
@@
Gaussian noise
We apply differential privacy to embedding computations by adding calibrated Gaussian noise. Specifically, for each query, noise sampled from $\mathcal{N}(0,\sigma^{2}I_{d})$ is added to returned embeddings, where $\sigma$ is calibrated to achieve $(\epsilon,\delta)$-differential privacy with $\epsilon=1.0$ and $\delta=10^{-5}$. Table 3 reports the robustness of UMID against a Differential Privacy (DP) defense in both the image and audio modalities. After injecting calibrated Gaussian noise ($\epsilon=1.0$), UMID exhibits only a moderate degradation: accuracy drops by roughly $6\%-12\%$, depending on the dataset and the number of samples. Concretely, for CLIP (ResNet-50x4) on CelebA, UMID still achieves $87.42\%$ accuracy with 1 sample and $85.95\%$ with 75 samples under the defense. Likewise, for CLAP on LibriSpeech, performance remains high at $87.88\%$ and $83.40\%$, and on CommonVoice it remains competitive at $75.90\%$ and $79.15\%$. Importantly, runtime is essentially unchanged across all settings (averaging $0.7$s), indicating that DP introduces negligible additional computational overhead for the attacker. Overall, these results suggest that while DP noise reduces the fidelity of embedding-based leakage, UMID’s reliance on geometric vulnerabilities remains effective under a moderate privacy budget; in practice, stronger privacy budgets ($\epsilon<1.0$) and/or complementary defenses may be necessary to fully secure these models.
Covert gibberish generation
In practice, a target model may incorporate input filters that flag overtly anomalous queries (e.g., nonsensical gibberish) and return misleading outputs, causing UMID to misclassify the presence of PII. To obtain more covert queries, we generate gibberish-like strings that remain visually and phonotactically plausible by replacing a small subset of characters with syllables drawn from a different language. For example, the detector can synthesize query texts by mixing English-style names with syllables from Arabic medical terminology. Concretely, we first prompt an LLM (e.g., GPT-3.5-turbo) to produce lists of common English initial and final syllables, then post-process the lists to remove duplicates and encourage coverage. We randomly concatenate syllables to form pseudo-English names (e.g., Karinix”, Zylogene”, “Renotyl”), and explicitly verify their novelty against a name lexicon to avoid collisions with real entities. Finally, we prompt the LLM to instantiate these syllable templates into full strings, yielding covert gibberish that closely resembles authentic names (see Table 6). Table 4 summarizes UMID’s robustness to the Covert Gibberish (CG) defense for both CLIP and CLAP. UMID remains highly resilient: CG reduces accuracy by only $1.23\%$$1.54\%$ across all datasets and training sample sizes. For example, on CelebA with 75 training samples, UMID still achieves $93.85\%$ accuracy under CG. Running time is also essentially unchanged (about 0.650.650.780.78s), suggesting that CG neither introduces meaningful computational overhead nor effectively disrupts the identification procedure. Overall, UMID preserves both effectiveness and efficiency under covert, text-like perturbations.
Table 4: UMID robustness evaluation under covert gibberish defense.
@@
6 Conclusion
@@ Appendix A Gibberish Samples
Table 5: Samples of randomly generated textual gibberish generated by GPT-3.5-turbo.
b+dh43u!f9de545w 53e(s3erg24=pnI<S fe3_5fg;@f34f.e2w/ d3|%542ferf34G_
4teh<E{434fef234ter 5gt%b@5435-hgF #4c3rd4scd324g ’2_:g435thyngb
Table 6: Covert gibberish that seem to be real names generated by GPT-3.5-turbo.
Karinix Zylogene Glycogenyx Zylotrax Vexilith Dynatrix
Exodynix Novylith Glycosyne Xenolynx Rynexis Delphylith
@@ Appendix B Theoretical Guarantees for Geometric Separation
In this appendix, we provide the detailed proofs for the theoretical claims made in Section 3.2. We establish that under mild geometric assumptions about the contrastive latent space, the proposed statistics $S_{n}$ and $D_{n}^{2}$ fundamentally separate member identities from non-members with high probability.
B.1 Preliminaries and Assumptions
@@
$\bar{v}_{n}$ $:=\frac{1}{n}\sum_{i=1}^{n}v^{(i)},\quad S_{n}:=v_{t}^{\top}\bar{v}_{n},\quad D_{n}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\|v^{(i)}-\bar{v}_{n}\|_{2}^{2}.$ (4)
@@
B.2 Population-Level Separation
@@
B.3 Finite-Sample Concentration

### 22 Apr 2026 v2

@@ IV-B Comparison with State-of-the-Art Methods
Figure 5: Empirical validation of geometric separation. (a) Convergence of similarity ($S_{n}$) and variability ($D_{n}^{2}$) as inversion runs ($n$) increase. (b) Cosine similarity distributions confirming the high-dimensional isotropy of non-members versus the directional bias of members.
IV-C Ablation Study
@@
IV-D Empirical Validation of Geometric Separation
To corroborate our theoretical guarantees (Theorem 3.1), we empirically validate the geometric behaviors of members and non-members in the latent space (Figure 5).
Convergence of Statistics ($S_{n}$ and $D_{n}^{2}$): Theoretically, members exhibit higher similarity ($S_{n}$) and lower variability ($D_{n}^{2}$) by concentrating on the memorized prototype. Empirically, as inversion runs ($n$) increase (Figure 5a), the variance of $S_{n}$ diminishes, and the member/non-member gap ($\Delta_{S}$) strictly stabilizes at $\approx 0.15$. Conversely, non-members’ $D_{n}^{2}$ converges to a significantly higher plateau, validating our dispersion bounds (Lemma B.5).
Verification of Non-member Isotropy: Assumption B.2 posits that non-members exhibit high-dimensional isotropy. We verify this by computing the cosine similarity between non-member embeddings ($v_{out}$) and random Gaussian vectors ($u$). As Figure 5b shows, this distribution strictly centers at 0 with minimal variance, confirming the isotropy condition. In contrast, members display a strong directional bias, demonstrating that $S_{n}$ and $D_{n}^{2}$ effectively capture fundamental geometric disparities rather than random noise.
V Robustness against Potential Defenses
@@
TABLE III: UMID robustness evaluation under differential privacy defense ($\epsilon=1.0$, $\delta=10^{-5}$).
@@
TABLE IV: UMID robustness evaluation under covert gibberish defense.
@@
Gaussian noise
We apply differential privacy to embedding computations by adding calibrated Gaussian noise. Specifically, for each query, noise sampled from $\mathcal{N}(0,\sigma^{2}I_{d})$ is added to returned embeddings, where $\sigma$ is calibrated to achieve $(\epsilon,\delta)$-differential privacy with $\epsilon=1.0$ and $\delta=10^{-5}$. Table III reports the robustness of UMID against a Differential Privacy (DP) defense in both the image and audio modalities. After injecting calibrated Gaussian noise ($\epsilon=1.0$), UMID exhibits only a moderate degradation: accuracy drops by roughly $6\%-12\%$, depending on the dataset and the number of samples. Concretely, for CLIP (ResNet-50x4) on CelebA, UMID still achieves $87.42\%$ accuracy with 1 sample and $85.95\%$ with 75 samples under the defense. Likewise, for CLAP on LibriSpeech, performance remains high at $87.88\%$ and $83.40\%$, and on CommonVoice it remains competitive at $75.90\%$ and $79.15\%$. Importantly, runtime is essentially unchanged across all settings (averaging $0.7$s), indicating that DP introduces negligible additional computational overhead for the attacker. Overall, these results suggest that while DP noise reduces the fidelity of embedding-based leakage, UMID’s reliance on geometric vulnerabilities remains effective under a moderate privacy budget; in practice, stronger privacy budgets ($\epsilon<1.0$) and/or complementary defenses may be necessary to fully secure these models.
Covert gibberish generation
In practice, a target model may incorporate input filters that flag overtly anomalous queries (e.g., nonsensical gibberish) and return misleading outputs, causing UMID to misclassify the presence of PII. To obtain more covert queries, we generate gibberish-like strings that remain visually and phonotactically plausible by replacing a small subset of characters with syllables drawn from a different language. For example, the detector can synthesize query texts by mixing English-style names with syllables from Arabic medical terminology. Concretely, we first prompt an LLM (e.g., GPT-3.5-turbo) to produce lists of common English initial and final syllables, then post-process the lists to remove duplicates and encourage coverage. We randomly concatenate syllables to form pseudo-English names (e.g., Karinix”, Zylogene”, “Renotyl”), and explicitly verify their novelty against a name lexicon to avoid collisions with real entities. Finally, we prompt the LLM to instantiate these syllable templates into full strings, yielding covert gibberish that closely resembles authentic names (see Table VI). Table IV summarizes UMID’s robustness to the Covert Gibberish (CG) defense for both CLIP and CLAP. UMID remains highly resilient: CG reduces accuracy by only $1.23\%$$1.54\%$ across all datasets and training sample sizes. For example, on CelebA with 75 training samples, UMID still achieves $93.85\%$ accuracy under CG. Running time is also essentially unchanged (about 0.650.650.780.78s), suggesting that CG neither introduces meaningful computational overhead nor effectively disrupts the identification procedure. Overall, UMID preserves both effectiveness and efficiency under covert, text-like perturbations.
VI Conclusion
@@
Author Contributions
Ruoxi Cheng contributed to the conceptualization and methodology, performed the preliminary experimental validation, and was responsible for writing the original draft. Yizhong Ding conducted the large-scale experimental validation and formal analysis. Jian Zhao and Tianle Zhang were responsible for the methodology refinement and validation. Hongyi Zhang and Haoxuan Ma contributed to the writing, reviewing, and editing of the manuscript. Yiyan Huang provided expertise in mathematical modeling and formal analysis, contributed to the writing, reviewing, and editing, and was responsible for funding acquisition and supervision. Xuelong Li provided conceptual guidance, project administration, and overall supervision. All authors have read and agreed to the published version of the manuscript.
@@ Appendix A Gibberish Samples
Table V and Table VI present examples of randomly generated gibberish and covert gibberish that mimics authentic names, respectively.
TABLE V: Samples of randomly generated textual gibberish generated by GPT-3.5-turbo.
b+dh43u!f9de545w 53e(s3erg24=pnI<S fe3_5fg;@f34f.e2w/
4teh<E{434fef234ter 5gt%b@5435-hgF #4c3rd4scd324g
TABLE VI: Covert gibberish that seem to be real names generated by GPT-3.5-turbo.
Karinix Zylogene Glycogenyx Zylotrax Vexilith
Exodynix Novylith Glycosyne Xenolynx Rynexis
@@ Appendix B Theoretical Guarantees for Geometric Separation
In this appendix, we provide the detailed proofs for the theoretical claims made in Section III-B. We establish that under mild geometric assumptions about the contrastive latent space, the proposed statistics $S_{n}$ and $D_{n}^{2}$ fundamentally separate member identities from non-members with high probability.
B-A Preliminaries and Assumptions
@@
$\bar{v}_{n}$ $:=\frac{1}{n}\sum_{i=1}^{n}v^{(i)},\quad S_{n}:=v_{t}^{\top}\bar{v}_{n},$
$D_{n}^{2}$ $:=\frac{1}{n}\sum_{i=1}^{n}\|v^{(i)}-\bar{v}_{n}\|_{2}^{2}.$ (4)
@@
B-B Population-Level Separation
@@
B-C Finite-Sample Concentration